OpenStudy (anonymous):

PLEASE HELP ME ! (: Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

1. point slope form is: y - y1 = m (x - x1) take points (3, -6) and plug them into the point slope form equation. take the slope (4) and plug it where it says x. try it by yourself, so you'll be able to do it on the test

OpenStudy (anonymous):

2. slope intercept form is y = mx + b 3. standard form is ax + by = c